OpenStudy (anonymous):

Finding critical points, after you find the zeros do I plug the test numbers into the derivative or the original function?

OpenStudy (anonymous):

you are better post the whole problem to make clear the question.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but i think you put the critical point into the original one to figure out the value of f and then compare them together

OpenStudy (anonymous):

don't forget the ending points. i mean interval

OpenStudy (anonymous):

f(x)=-1/x^2 on the interval from .5<=x,=2 f '(x) =-x^2x^-3 then set to 0 ssolve for x, x=0. so I know I plug in the end points and x values to find max/min. But do I plug into the derivative or the original function?

OpenStudy (anonymous):

the interval is not clear, \[0.5\le x \le 2\]is it right?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

and your f'(x) must be = 2/x^3. tell me if i am wrong. I 'm dealing with critical point, too. we can discuss and help each other

OpenStudy (anonymous):

ok great! I think its 2x^-3 though

OpenStudy (anonymous):

or 2 1/x^3

OpenStudy (anonymous):

why? your f(x) = -x^-2. take derivative of it, ok, you are right now, not previous result

OpenStudy (anonymous):

so, you don't have critical point except ending point. because x cannot be 0 to make f'x defined

OpenStudy (anonymous):

oh that makes sense ,

OpenStudy (anonymous):

count f(0.5) and f(2) and compare them only

OpenStudy (anonymous):

So then x wont have a 0 then because its in a denominator..... ok I see so then that means that i doesn't cross th x axis then right?

OpenStudy (anonymous):

it doesn't mean that. the critical points talk about max, min of function

OpenStudy (anonymous):

But then do you use f(.5) or f ' (.5) ?

OpenStudy (anonymous):

f(0.5)= -2 is local min and f(2)=-1/4 is local max

OpenStudy (anonymous):

oh right this tells me when the tan line is =0,

OpenStudy (anonymous):

if consider of the value on interval [0.5,2] they are absolute max and min

OpenStudy (anonymous):

not crossing the axis

OpenStudy (anonymous):

i think you confuse. the critical point tells us about the tangent line is horizontal line with the slope =0. in your case, you don't have it. if you have it is vertical line with x =0

OpenStudy (anonymous):

that is y-axis

OpenStudy (anonymous):

in other word, it has y axis as its asymtote

OpenStudy (anonymous):

verify by drawing the graph, trivially see that

OpenStudy (anonymous):

Hello, do I lost you?

OpenStudy (anonymous):

i'm back sorry

OpenStudy (anonymous):

so i plot my intervals on line, then plug test numbers in to each interval right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so f'(-1)=-2 to the left of .5,dec. f'(1)=2 in between .5 and 2 inc and f'(3)=2/27 to rt of 2 = inc

OpenStudy (anonymous):

why you put those points into f'?. just 2 ending points into ORIGINAL function

OpenStudy (anonymous):

so i should put the end points into the original to test?

OpenStudy (anonymous):

not to test, to figure out which point is maximum, or minimum value of f

OpenStudy (anonymous):

? the first derivative test to see if inc/dec right

OpenStudy (anonymous):

yes, of course. if you have f(0.5)> f(2). does it mean your function decreases?

OpenStudy (anonymous):

so after the first derive test then you put the x's into original to find the corresponding y values?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i see i'm gonna try another one and get back to you thanks

OpenStudy (anonymous):

the derivative just to make sure that among the interval, your function has another up or down or not. if it has another critical points, put those points to ORIGINAL one to know it it's up or down

OpenStudy (anonymous):

no critical point, f decreases from a, to b.